OpenStudy (samara8954):

Seraphina is driving two hours to visit her family. For the first hour, she traveled at a speed of 63 miles per hour. Then, in the second hour, she traveled at a speed of 66 miles per hour. What is the percentage increase of Seraphina's speed? If necessary, round to the nearest tenth of a percent. 4.8% 95.2% 4.5% 95.5%

OpenStudy (samara8954):

@dtan5457

OpenStudy (dtan5457):

Same thing as before. Subtract new value from old 66-63

OpenStudy (dtan5457):

3 divided by original value of 63 is>?

OpenStudy (calculusxy):

63 of ___ = 3 63 => the original price ___ => the percent (which you are trying to find) 3 => the amount in which the price was increased by

OpenStudy (samara8954):

@dtan5457

OpenStudy (dtan5457):

You went up 3 mph.. Basically your trying to find out how much faster or what percentage 3 is of 63<<the original speed

OpenStudy (dtan5457):

therefore 3/63=?

OpenStudy (samara8954):

but the answer isn't on thier @dtan5457

OpenStudy (calculusxy):

To solve for the blank, you use this equation. \[\frac{ is }{ of }=\frac{ percent }{ 100 }\] Remember that "is" is substituted for the value after the equal sign (in this case it is 3). "of" is substituted for the original value (which is 63) \[\frac{ 3 }{ 63 }=\frac{ x }{ 100 }\] You can solve for x by cross-multiplying. \[3 \times 100 = 300\] \[63 \times x = 63x\] So you have 300 = 63x. Divide 300 by 63 to get the quotient. And then move the decimal place two places to the right to get your answer.

OpenStudy (dtan5457):

The answer IS on there, you just have to round the decimals.

OpenStudy (dtan5457):

0.04761904761 Round that..^

OpenStudy (samara8954):

0.48

OpenStudy (calculusxy):

Wait. But you move the decimal place two places to the right and then round to the nearest tenths.

OpenStudy (samara8954):

so 4.8 is the answer

OpenStudy (calculusxy):

Since you are converting from a decimal to a percent.

OpenStudy (calculusxy):

Yes

OpenStudy (dtan5457):

correct @Samara8954